Grade 12General Physics

Explain about Prism:

outpass2
10 Years agoGrade 12

2 Answers

OUTPASS
ApprovedApproved Tutor Answer10 Years ago
a) Angle of the prism A = r1 + r2 and Angle of deviation (d). =I + e – A
I – angle of incidence,
e – angle of emergence,
A – angle of prism,
r1 – angle of refraction at first retracing face,
r2 – angle of refraction at second refracting face.
A + Dm 
sin 
b) μ = 
sin A
2
 
Surendra
10 Years ago
) Angle of the prism A = r1 + r2 and Angle of deviation (d). =I + e – AI – angle of incidence,e – angle of emergence,A – angle of prism,r1 – angle of refraction at first retracing face,r2 – angle of refraction at second refracting face.A + Dm sin b) μ = sin A2